OpenStudy (firejay5):

Write an equation for the ellipse that satisfies the conditions: I am not sure to begin honestly; I just need a good starting point to begin. 8. endpoints of major axis (-6, 0) and (6, 0) and foci ( -sqrt 32, 0) and ( sqrt 32, 0)

OpenStudy (firejay5):

@ParthKohli

ganeshie8 (ganeshie8):

`endpoints of major axis (-6, 0) and (6, 0)` what does this tell us about the value of semi major axis length : `a` ?

OpenStudy (firejay5):

It tells us that it goes up 6 and down -6

ganeshie8 (ganeshie8):

not exactly, ` (-6, 0) and (6, 0)` are points on x axis, right ?

OpenStudy (firejay5):

left and right

ganeshie8 (ganeshie8):

yes that means \(\large a = 6\)

ganeshie8 (ganeshie8):

\(\large c = \sqrt{32}\)

ganeshie8 (ganeshie8):

find the value of \(\large b\) using below relation : \[\large c^2 = a^2 - b^2\]

OpenStudy (firejay5):

is that right b = +/- 2 @ganeshie8

ganeshie8 (ganeshie8):

yes, since b is a length it can only by positive so b = +2

ganeshie8 (ganeshie8):

can you write the equation of ellipse now ?

ganeshie8 (ganeshie8):

you have \(\large a = 6\) and \(\large b = 2\) equation of ellipse would be ?

OpenStudy (firejay5):

\[\frac{ x^2 }{ 36 } + \frac{ y^2 }{ 4 }\]

OpenStudy (firejay5):

= 1

ganeshie8 (ganeshie8):

Excellent !

OpenStudy (firejay5):

is that right?

ganeshie8 (ganeshie8):

Yep ! you may check with wolfram when in doubt http://www.wolframalpha.com/input/?i=focus+x%5E2%2F36%2By%5E2%2F4%3D1 scroll down

OpenStudy (firejay5):

what if it was switched like major axis was (0, 12) and (0, -12) and foci was at (0, sqrt 23) and (0, - sqrt 23) would the problem working be the same as the one we just did, but different answers.

OpenStudy (firejay5):

@ganeshie8

ganeshie8 (ganeshie8):

yes it will be more or less same, but the most important thing is to find out whether the ellipse is horizontal or vertical

ganeshie8 (ganeshie8):

`major axis was (0, 12) and (0, -12) ` since the major axis is on y axis, the ellipse is VERTICAL and \(\large a = 12\)

ganeshie8 (ganeshie8):

\(\large c = \sqrt{23}\) as usual find the value of \(\large b \)

ganeshie8 (ganeshie8):

but be careful to use the correct equation for `vertical ellipses` : \[\large \dfrac{x^2}{b^2} + \dfrac{y^2}{a^2} = 1\]

ganeshie8 (ganeshie8):

\(\large a\) goes below the `y` variable ^^

OpenStudy (firejay5):

its horizontal \[\frac{ x^2 }{ 144 } + \frac{ y^2 }{ 11 } = 1\]

ganeshie8 (ganeshie8):

nope, it has to be vertical because you're given : `major axis was (0, 12) and (0, -12) ` major axis is along y axis, right ?

OpenStudy (firejay5):

well switch 144 and 11 then

OpenStudy (firejay5):

11 is 121

ganeshie8 (ganeshie8):

a = 12, b = 11 yes ! the equation would be \[\large \frac{ x^2 }{ 121 } + \frac{ y^2 }{ 144 } = 1\]

OpenStudy (firejay5):

so I had it right just messed up on denominator

ganeshie8 (ganeshie8):

yes refer to that chart, its very useful